What is the dimensional formula of gravitational force

Grade:11

4 Answers

Arun
6 years ago

Universal Constant of Gravitation is represented by G and is derived from Newton’s law of gravitation. Newton’s law states that
F=(Gm1m2)/r2
So
G=(Fr2)/m1m2

Where F is Force between masses m1, m2 at a distance r

Dimensional Formula of Force= M1L1T-2

Thus
Dimensional Formula of Universal Constant of Gravitation= (M1L1T-2 x L2)/(Mx M1)

Dimensional Formula of Universal Constant of Gravitation = M-1L3T-2
SI unit of Universal Constant of Gravitation is Nm2kg-2

or

Dimensional Formula of Gravitational constant = M-1L3T-2

anjace
6 years ago
g=force*(radius)^2/m1*m2
where m1 and m2 are masses of earth and object respectively.
Anjali kumari
5 years ago
It's dimension will be MLT^-1. We do not need to care about which force is. For gravitational force or magnetic force the dimension will be same as the dimension of force.
 
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Universal Constant of Gravitation is represented by G and is derived from Newton’s law of gravitation. Newton’s law states that
F=(Gm1m2)/r2
So
G=(Fr2)/m1m2
Where F is Force between masses m1, m2at a distance rDimensional Formula of Force= M1L1T-2Thus
Dimensional Formula of Universal Constant of Gravitation= (M1L1T-2x L2)/(M1x M1)
Dimensional Formula of Universal Constant of Gravitation = M-1L3T-2
SI unit ofUniversal Constant of Gravitation is Nm2kg-2
or
Dimensional Formula of Gravitational constant = M-1L3T-2

Thanks and Regards